1. Suppose the investor finds a tax free security that pays 6%. The investors tax rate on taxable securities is 30%. Find the tax equivalent yield. 2. Find the yield on a 30 day debt security if Rn-2%, DP -3%, LP 2% and TA-. 1 %. 3. Assume tl,--6%,-3%, ti,-2%. Find the one and two year forward rates for a one year security.
Answer #1

1.

Tax equivalent yield is used to compare the yields of tax exempt securities to that of taxable securities. It is given by the formula:

yield on tar free security 1-tar rate Tar equivalent yield

0.06 Tax equivalent yield 1-0.3

= 8.57%

--------------------------------------------------------------

2.

The formula to calculate yield of an n day security is:

n- A

where,

  • Y_{n} is the yield to be calculated
  • 71 is the risk free yield
  • DP is the default premium
  • LP is the liquidity premium
  • TA is adjustments due to difference in tax status

Yield = 2% + 3% + 2% + 1% = 8%

3.

Forward rates are interest rates that are decided now, for a transaction that is going to happen sometime in the future. The figure below shows the spot rates rates for periods 1, 2 and 3 respectively as given in the question. The one year forward rate is then denoted by f_{_{1}}. The rate is fixed today between investor and borrower i.e. at time period 0. The actual transaction happens at period 1 and the transaction concludes at period 2.

f2 0 2% 1 3 3% 6%

Now to calculate one year forward rate f_{_{1}} we can use the following equation:

(1+0.02)(1+f_{1})=(1+0.03)^{2}

this equation simply means that the return an investor would get by holding a 1 year security and a 1 year forward security is equal to him holding a 2 year security i.e. there is no opportunity for arbitrage.

On solving we get

f_{1}=4%

Similarly we can write an equation for 2 year forward rate as:

(1 + 0,03 )2(1 + /2) = (1 + 0,06)3

On solving, we get

f_{2}=12.3%